# Stellar Superradiance and Low-Energy Absorption in Dense Nuclear Media

### v1

@@ 2.1 Phase Space Integral
$\Gamma^{A}=\frac{1}{2\omega_{b}}\left[\prod^{4}_{i=1}\int\frac{{\rm d}^{3}\vec{p}_{i}}{2\omega_{i}(2\pi)^{3}}\right]\,f_{1}f_{2}(1-f_{3})(1-f_{4})\,\delta^{4}(p_{1}+p_{2}-p_{3}-p_{4}+p_{b})\,(2\pi)^{4}\,\frac{\sum_{s}|\mathcal{M}_{\mathrm{SC}}|^{2}}{4}.$ (S19)
@@
$\begin{split}&\Gamma^{A}_{a}\simeq\frac{p_{F}^{4}}{2^{15}\pi^{8}\omega_{b}}\frac{m_{b}^{2}}{\Gamma_{\rm col}^{2}}\frac{g_{\pi N}^{4}m_{\psi}^{2}\,g_{a}^{2}}{m_{\pi}^{4}}\left[\frac{256}{3}\frac{|\vec{q}_{b}|^{2}}{m_{b}^{2}}+128\frac{p_{F}^{2}}{m_{\psi}^{2}}\right]A^{\rm SS}B^{\rm SS},\\ &\Gamma^{A}_{\rm EDM}\simeq\frac{p_{F}^{4}}{2^{15}\pi^{8}\omega_{b}}\frac{m_{b}^{2}}{\Gamma_{\rm col}^{2}}\frac{g_{\pi N}^{4}m_{\psi}^{2}\,g_{\rm EDM}^{2}}{m_{\pi}^{4}}\frac{256}{3}A^{\rm SS}B^{\rm SS},\\ &\Gamma^{A}_{\rm MDM}\simeq\frac{p_{F}^{4}}{2^{15}\pi^{8}\omega_{b}}\frac{m_{b}^{2}}{\Gamma_{\rm col}^{2}}\frac{g_{\pi N}^{4}m_{\psi}^{2}\,g_{\rm MDM}^{2}}{m_{\pi}^{4}}64\left[\frac{|\vec{q}_{b}|^{2}}{m_{b}^{2}}A^{\rm SS}_{1}+\frac{p_{F}^{2}}{m_{\psi}^{2}}A^{\rm SS}_{2}\right]B^{\rm SS}.\end{split}$ (S20)
@@
$\begin{split}B^{\rm SS}&=T^{4}\left[\prod_{i=1}^{4}\int_{-\mu_{i}/T}^{\infty}{\rm d}y_{i}\right]\,\delta(\omega_{1}+\omega_{2}-\omega_{3}-\omega_{4}+\omega_{b})\,f_{1}f_{2}(1-f_{3})(1-f_{4})\\ &\approx T^{3}\left[\prod_{i=1}^{4}\int_{-\infty}^{\infty}{\rm d}y_{i}f_{i}\right]\delta\bigg(\sum_{i}y_{i}+y_{b}\bigg)\\ &\approx\frac{T^{3}}{6}\frac{y_{b}(y_{b}^{2}+4\pi^{2})}{1-e^{-y_{b}}}\\ &\simeq\frac{2\pi^{2}}{3}T^{3},\end{split}$ (S21)
@@
$A^{\rm SS}=$ $\left[\prod_{i=1}^{4}\int{\rm d}\Omega_{i}\right]\,\delta^{3}(\vec{p}_{1}+\vec{p}_{2}-\vec{p}_{3}-\vec{p}_{4})\left[\left\{\mathcal{M}_{k}^{2}\right\}+\left\{\mathcal{M}_{l}^{2}\right\}+\left\{\mathcal{M}_{kl}^{2}\right\}\right],$ (S22)
$A^{\rm SS}_{1}=$ $\left[\prod_{i=1}^{4}\int{\rm d}\Omega_{i}\right]\,\delta^{3}(\vec{p}_{1}+\vec{p}_{2}-\vec{p}_{3}-\vec{p}_{4})\left[\frac{4}{3}\{\mathcal{M}_{k}^{2}\}+\frac{4}{3}\{\mathcal{M}_{l}^{2}\}+\frac{8}{9}\{(\mathcal{M}_{kl}^{2}\}\right],$
@@
$\begin{split}A^{\rm SS}=&\frac{1}{p_{F}^{8}}\left[\prod_{i=1}^{4}\int{\rm d}^{3}\vec{p}_{i}\right]\delta^{3}\left(\vec{p}_{1}+\vec{p}_{2}-\vec{p}_{3}-\vec{p}_{4}\right)\prod_{i=1}^{4}\delta(|\vec{p}_{i}|-p_{F})\left[\left\{\mathcal{M}_{k}^{2}\right\}+\left\{\mathcal{M}_{l}^{2}\right\}+\left\{\mathcal{M}_{kl}^{2}\right\}\right],\\ =&\frac{1}{p_{F}^{8}}\int{\rm d}^{3}\vec{p}_{1}\,{\rm d}^{3}\vec{k}\,{\rm d}^{3}\vec{l}\,\,\delta\left(|\vec{p}_{1}|-p_{F}\right)\delta\left(|\vec{p}_{1}+\vec{k}+\vec{l}|-p_{F}\right)\delta\left(|\vec{p}_{1}+\vec{k}|-p_{F}\right)\delta\left(|\vec{p}_{1}+\vec{l}|-p_{F}\right)\\ &\times\left[\left\{\mathcal{M}_{k}^{2}\right\}+\left\{\mathcal{M}_{l}^{2}\right\}+\left\{\mathcal{M}_{kl}^{2}\right\}\right]\\ \simeq&16\pi^{2}p_{F}^{-4}\int_{0}^{2p_{F}}{\rm d}|\vec{k}|\times 2\int_{0}^{\sqrt{4p_{F}^{2}-|\vec{k}|^{2}}}{\rm d}|\vec{l}|\,\frac{\left\{\mathcal{M}_{k}^{2}\right\}+\left\{\mathcal{M}_{l}^{2}\right\}+\left\{\mathcal{M}_{kl}^{2}\right\}}{\sqrt{4p_{F}^{2}-|\vec{k}|^{2}-|\vec{l}|^{2}}}\\ =&32\pi^{3}p^{-3}_{F}F(\xi).\end{split}$ (S23)
where $F(\xi)\approx 1.67$ with $\xi=m_{\pi}/(2p_{F})$, following the definition in Eq. (S11). We have also used the approximate orthogonality $\vec{k}\cdot\vec{l}\simeq 0$.
@@
$\begin{split}F_{\rm MDM}^{1}(\xi)&=\frac{2}{9}\left[\pi\bigg(\frac{2\xi}{\sqrt{2\xi^{2}+1}}-13\bigg)\xi-\frac{6}{\xi^{2}+1}-\frac{4\xi^{2}}{\sqrt{1+2\xi^{2}}}\mathrm{arctan}(\frac{\xi^{2}}{\sqrt{1+2\xi^{2}}})+26\xi\arctan{\xi}+22\right]\\ &\approx 2.01,\\ F_{\rm MDM}^{2}(\xi)&=\frac{2}{9(1+\xi^{2})}\left[20-105\xi^{4}-61\xi^{2}+3(\xi^{2}+1)(35\xi^{2}-3)\xi\arctan{\xi}\right]\approx 3.36.\end{split}$ (S25)
@@
${\Gamma}^{\rm A}_{a}\simeq\frac{g_{\pi N}^{4}m_{\psi}^{2}g_{a}^{2}}{m_{\pi}^{4}}\frac{m_{b}^{2}}{\Gamma_{\rm col}^{2}}\left[\frac{256}{3}\frac{|\vec{q}_{b}|^{2}}{m_{b}^{2}}+128\frac{p_{F}^{2}}{m_{\psi}^{2}}\right]\frac{p_{F}T^{2}}{1536\pi^{3}\omega_{b}/T}F(\xi),$ (S26)
@@
2.2 Stellar Superradiance Rates
@@
$\Gamma^{\rm SR}=\frac{\omega_{b}}{T}\frac{\Omega_{S}m-\omega_{b}}{\omega_{b}}\int\phi^{*}\Gamma_{A}\phi\,{\rm d}^{3}\vec{r},$ (S27)
@@
$C\equiv p_{F}T^{2}\frac{g_{\pi N}^{4}m_{\psi}^{2}}{m_{\pi}^{4}}\frac{\Omega_{S}m-\omega_{b}}{\omega_{b}}\bigg(\frac{R_{S}}{r_{b}}\bigg)^{3}.$ (S29)
3 Effective Action Approach to Boson Damping in a Fermion Medium
@@
$\mathcal{L}_{\text{int}}\;=\;j_{A}(x)\,\mathcal{O}^{A}[\Phi](x),$ (S30)
@@
$\bigl(\Box+m_{b}^{2}\bigr)\,\Phi(x)\;+\;\Bigl\langle\frac{\partial\mathcal{L}_{\text{int}}}{\partial\Phi(x)}\Bigr\rangle\;=\;0,$ (S31)
@@
$\Gamma[\Phi]\;=\;\frac{1}{2}\int{\rm d}^{4}x\,{\rm d}^{4}y\;\mathcal{O}^{A}[\Phi](x)\,G^{R}_{AB}(x-y)\,\mathcal{O}^{B}[\Phi](y),$ (S32)
@@
$G^{R}_{AB}(x-y)\;=\;-\,i\,\Theta(x^{0}-y^{0})\,\bigl\langle[\,j_{A}(x),\,j_{B}(y)\,]\bigr\rangle$ (S33)
@@
$\Gamma[\Phi]\;\simeq\;\frac{1}{2}\int{\rm d}^{4}X\;\mathcal{O}^{A}[\Phi](X)\,\mathcal{O}^{B}[\Phi](X)\,G_{AB}(\omega_{b}),$ (S34)
@@
$\text{Im}\,G_{AB}(\omega_{b})=-\pi\sum_{i,f}\rho_{i}\,(2\pi)^{3}\delta^{3}(\vec{p}_{f}-\vec{p}_{i})\left[\delta(\omega_{b}+E_{f}-E_{i})\,M^{AB}_{if}-\delta(\omega_{b}-E_{f}+E_{i})\,M^{AB}_{fi}\right],$ (S36)
@@
Substituting Eq. (S36) into the effective action (S34) and varying then gives the imaginary part of the interaction term in Eq. (S31).
@@
$\frac{\delta\,\text{Im}\Gamma}{\delta\Phi(x)}=-\partial_{i}\!\left[\text{Im}G_{jj}(\omega_{b})\,\partial^{i}\Phi(x)\right]=-\text{Im}G_{jj}(\omega_{b})\,\nabla^{2}\Phi(x).$ (S37)
@@
By converting the sums in Eq. (S36) into phase-space integrals, the ensemble weights $\rho_{i}$ give rise to the thermal occupation factors, such as $f_{1}f_{2}(1-f_{3})(1-f_{4})$ in Eq. (3.7). The factor $\text{Im}\,G_{jj}(\omega_{b})\,\nabla^{2}$ in Eq. (S37) can then be identified with the term $(\Gamma^{A}-\Gamma^{E})\,\partial_{t}$ in Eq. (3.1). In this correspondence, the spatial momentum factors that appear in the squared matrix elements are replaced by spatial gradient operators acting on $\Phi$ in position space.

### v2

@@ B.1 Phase Space Integral
$\Gamma^{A}=\frac{1}{2\omega_{b}}\left[\prod^{4}_{i=1}\int\frac{{\rm d}^{3}\vec{p}_{i}}{2\omega_{i}(2\pi)^{3}}\right]\,f_{1}f_{2}(1-f_{3})(1-f_{4})\,\delta^{4}(p_{1}+p_{2}-p_{3}-p_{4}+p_{b})\,(2\pi)^{4}\,\frac{\sum_{s}|\mathcal{M}_{\mathrm{SC}}|^{2}}{4}.$ (B.8)
@@
$\begin{split}&\Gamma^{A}_{a}\simeq\frac{p_{F}^{4}}{2^{15}\pi^{8}\omega_{b}}\frac{m_{b}^{2}}{\Gamma_{\rm col}^{2}}\frac{g_{\pi N}^{4}m_{\psi}^{2}\,g_{a}^{2}}{m_{\pi}^{4}}\left[\frac{256}{3}\frac{|\vec{q}_{b}|^{2}}{m_{b}^{2}}+128\frac{p_{F}^{2}}{m_{\psi}^{2}}\right]A^{\rm SS}B^{\rm SS},\\ &\Gamma^{A}_{\rm EDM}\simeq\frac{p_{F}^{4}}{2^{15}\pi^{8}\omega_{b}}\frac{m_{b}^{2}}{\Gamma_{\rm col}^{2}}\frac{g_{\pi N}^{4}m_{\psi}^{2}\,g_{\rm EDM}^{2}}{m_{\pi}^{4}}\frac{256}{3}A^{\rm SS}B^{\rm SS},\\ &\Gamma^{A}_{\rm MDM}\simeq\frac{p_{F}^{4}}{2^{15}\pi^{8}\omega_{b}}\frac{m_{b}^{2}}{\Gamma_{\rm col}^{2}}\frac{g_{\pi N}^{4}m_{\psi}^{2}\,g_{\rm MDM}^{2}}{m_{\pi}^{4}}64\left[\frac{|\vec{q}_{b}|^{2}}{m_{b}^{2}}A^{\rm SS}_{1}+\frac{p_{F}^{2}}{m_{\psi}^{2}}A^{\rm SS}_{2}\right]B^{\rm SS}.\end{split}$ (B.9)
@@
$\begin{split}B^{\rm SS}&=T^{4}\left[\prod_{i=1}^{4}\int_{-\mu_{i}/T}^{\infty}{\rm d}y_{i}\right]\,\delta(\omega_{1}+\omega_{2}-\omega_{3}-\omega_{4}+\omega_{b})\,f_{1}f_{2}(1-f_{3})(1-f_{4})\\ &\approx T^{3}\left[\prod_{i=1}^{4}\int_{-\infty}^{\infty}{\rm d}y_{i}f_{i}\right]\delta\bigg(\sum_{i}y_{i}+y_{b}\bigg)\\ &\approx\frac{T^{3}}{6}\frac{y_{b}(y_{b}^{2}+4\pi^{2})}{1-e^{-y_{b}}}\\ &\simeq\frac{2\pi^{2}}{3}T^{3},\end{split}$ (B.10)
@@
$A^{\rm SS}=$ $\left[\prod_{i=1}^{4}\int{\rm d}\Omega_{i}\right]\,\delta^{3}(\vec{p}_{1}+\vec{p}_{2}-\vec{p}_{3}-\vec{p}_{4})\left[\left\{\mathcal{M}_{k}^{2}\right\}+\left\{\mathcal{M}_{l}^{2}\right\}+\left\{\mathcal{M}_{kl}^{2}\right\}\right],$ (B.11)
$A^{\rm SS}_{1}=$ $\left[\prod_{i=1}^{4}\int{\rm d}\Omega_{i}\right]\,\delta^{3}(\vec{p}_{1}+\vec{p}_{2}-\vec{p}_{3}-\vec{p}_{4})\left[\frac{4}{3}\{\mathcal{M}_{k}^{2}\}+\frac{4}{3}\{\mathcal{M}_{l}^{2}\}+\frac{8}{9}\{\mathcal{M}_{kl}^{2}\}\right],$
@@
$\begin{split}A^{\rm SS}=&\frac{1}{p_{F}^{8}}\left[\prod_{i=1}^{4}\int{\rm d}^{3}\vec{p}_{i}\right]\delta^{3}\left(\vec{p}_{1}+\vec{p}_{2}-\vec{p}_{3}-\vec{p}_{4}\right)\prod_{i=1}^{4}\delta(|\vec{p}_{i}|-p_{F})\left[\left\{\mathcal{M}_{k}^{2}\right\}+\left\{\mathcal{M}_{l}^{2}\right\}+\left\{\mathcal{M}_{kl}^{2}\right\}\right],\\ =&\frac{1}{p_{F}^{8}}\int{\rm d}^{3}\vec{p}_{1}\,{\rm d}^{3}\vec{k}\,{\rm d}^{3}\vec{l}\,\,\delta\left(|\vec{p}_{1}|-p_{F}\right)\delta\left(|\vec{p}_{1}+\vec{k}+\vec{l}|-p_{F}\right)\delta\left(|\vec{p}_{1}+\vec{k}|-p_{F}\right)\delta\left(|\vec{p}_{1}+\vec{l}|-p_{F}\right)\\ &\times\left[\left\{\mathcal{M}_{k}^{2}\right\}+\left\{\mathcal{M}_{l}^{2}\right\}+\left\{\mathcal{M}_{kl}^{2}\right\}\right]\\ \simeq&16\pi^{2}p_{F}^{-4}\int_{0}^{2p_{F}}{\rm d}|\vec{k}|\times 2\int_{0}^{\sqrt{4p_{F}^{2}-|\vec{k}|^{2}}}{\rm d}|\vec{l}|\,\frac{\left\{\mathcal{M}_{k}^{2}\right\}+\left\{\mathcal{M}_{l}^{2}\right\}+\left\{\mathcal{M}_{kl}^{2}\right\}}{\sqrt{4p_{F}^{2}-|\vec{k}|^{2}-|\vec{l}|^{2}}}\\ =&32\pi^{3}p^{-3}_{F}F(\xi),\end{split}$ (B.12)
where $F(\xi)\approx 1.67$ with $\xi=m_{\pi}/(2p_{F})$, following the definition in Eq. (A.9). We have also used the approximate orthogonality $\vec{k}\cdot\vec{l}\simeq 0$.
@@
$\begin{split}F_{\rm MDM}^{1}(\xi)&=\frac{2}{9}\left[\pi\bigg(\frac{2\xi}{\sqrt{2\xi^{2}+1}}-13\bigg)\xi-\frac{6}{\xi^{2}+1}-\frac{4\xi^{2}}{\sqrt{1+2\xi^{2}}}\mathrm{arctan}(\frac{\xi^{2}}{\sqrt{1+2\xi^{2}}})+26\xi\arctan{\xi}+22\right]\\ &\approx 2.01,\\ F_{\rm MDM}^{2}(\xi)&=\frac{2}{9(1+\xi^{2})}\left[20-105\xi^{4}-61\xi^{2}+3(\xi^{2}+1)(35\xi^{2}-3)\xi\arctan{\xi}\right]\approx 3.36.\end{split}$ (B.14)
@@
${\Gamma}^{\rm A}_{a}\simeq\frac{g_{\pi N}^{4}m_{\psi}^{2}g_{a}^{2}}{m_{\pi}^{4}}\frac{m_{b}^{2}}{\Gamma_{\rm col}^{2}}\left[\frac{256}{3}\frac{|\vec{q}_{b}|^{2}}{m_{b}^{2}}+128\frac{p_{F}^{2}}{m_{\psi}^{2}}\right]\frac{p_{F}T^{2}}{1536\pi^{3}\omega_{b}/T}F(\xi),$ (B.15)
@@
B.2 Stellar Superradiance Rates
@@
$\Gamma^{\rm SR}=\frac{\omega_{b}}{T}\frac{\Omega_{S}m-\omega_{b}}{\omega_{b}}\int\phi^{*}\Gamma_{A}\phi\,{\rm d}^{3}\vec{r},$ (B.16)
@@
$C\equiv p_{F}T^{2}\frac{g_{\pi N}^{4}m_{\psi}^{2}}{m_{\pi}^{4}}\frac{\Omega_{S}m-\omega_{b}}{\omega_{b}}\bigg(\frac{R_{S}}{r_{b}}\bigg)^{3}.$ (B.18)
Appendix C Effective Action Approach to Boson Damping in a Fermion Medium
@@
$\mathcal{L}_{\text{int}}\;=\;j_{A}(x)\,\mathcal{O}^{A}[\Phi](x),$ (C.1)
@@
$\bigl(\Box+m_{b}^{2}\bigr)\,\Phi(x)\;+\;\Bigl\langle\frac{\partial\mathcal{L}_{\text{int}}}{\partial\Phi(x)}\Bigr\rangle\;=\;0,$ (C.2)
@@
$\Gamma[\Phi]\;=\;\frac{1}{2}\int{\rm d}^{4}x\,{\rm d}^{4}y\;\mathcal{O}^{A}[\Phi](x)\,G^{R}_{AB}(x-y)\,\mathcal{O}^{B}[\Phi](y),$ (C.3)
@@
$G^{R}_{AB}(x-y)\;=\;-\,i\,\Theta(x^{0}-y^{0})\,\bigl\langle[\,j_{A}(x),\,j_{B}(y)\,]\bigr\rangle$ (C.4)
@@
$\Gamma[\Phi]\;\simeq\;\frac{1}{2}\int{\rm d}^{4}X\;\mathcal{O}^{A}[\Phi](X)\,\mathcal{O}^{B}[\Phi](X)\,G_{AB}(\omega_{b}),$ (C.5)
@@
$\text{Im}\,G_{AB}(\omega_{b})=-\pi\sum_{i,f}\rho_{i}\,(2\pi)^{3}\delta^{3}(\vec{p}_{f}-\vec{p}_{i})\left[\delta(\omega_{b}+E_{f}-E_{i})\,M^{AB}_{if}-\delta(\omega_{b}-E_{f}+E_{i})\,M^{AB}_{fi}\right],$ (C.7)
@@
Substituting Eq. (C.7) into the effective action (C.5) and varying then gives the imaginary part of the interaction term in Eq. (C.2).
@@
$\frac{\delta\,\text{Im}\Gamma}{\delta\Phi(x)}=-\partial_{i}\!\left[\text{Im}G_{jj}(\omega_{b})\,\partial^{i}\Phi(x)\right]=-\text{Im}G_{jj}(\omega_{b})\,\nabla^{2}\Phi(x).$ (C.8)
@@
By converting the sums in Eq. (C.7) into phase-space integrals, the ensemble weights $\rho_{i}$ give rise to the thermal occupation factors, such as $f_{1}f_{2}(1-f_{3})(1-f_{4})$ in Eq. (3.7). The factor $\text{Im}\,G_{jj}(\omega_{b})\,\nabla^{2}$ in Eq. (C.8) can then be identified with the term $(\Gamma^{A}-\Gamma^{E})\,\partial_{t}$ in Eq. (3.1). In this correspondence, the spatial momentum factors that appear in the squared matrix elements are replaced by spatial gradient operators acting on $\Phi$ in position space.